OpenStudy (anonymous):

Factor the GCF and then the resulting trinomial: 3x3 - 3x2 - 6x A. x(3x - 6)(x + 1) B. 3(x - 2)(x + 1) C. 3x(x - 2)(x + 1) D. prime

OpenStudy (apoorvk):

C.

OpenStudy (arnavguddu):

3x3-3x2-6x = 3x (x2-x-2) = 3x(x2-2x+x-2)=3x(x(x-2)+1(x-2)) = 3x(x-2)(x+1) Ans-C